Prev Next

SW&IT Meeting Minutes 2022-07-11

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • xxx
  • This week
    • xxx

YZ

  • Previous:
    • TMS related work.
    • LBC code study and planning.
    • Some MODS new utility discussion with Olga.
    • Some other misc support stuff.
  • Next:
    • Mainly focus on TMS and LBC work.
    • Tentative TMS test (Wed or Fri).
    • Working on some new MODS utilities.
    • Other possible misc support stuff.

DE

  • Last week:
    • Veeam sanity check quote, hard no from us. We'll just have to juggle the mountain VMs, not a big deal.
    • PG soft offboarding, seems happy with the new level of LBT mail
    • Stripped LBTI stuff out of our Google account, will revisit SSD 2023
    • RR Vault issues, was mostly a matter of organization. I cleaned old versions out.
    • Discussion with SE re: new scratch/buffer disk for NOMIC computer. SE is good with the more fragile quad-bit SSD, so they'll move forward with that
    • No freakout emails from Barry, so the UITS work didn't impact ops.
  • This week:
    • SSD stuff, will nail down the best date and go in to do the Tucson arrays and work out when I can have some of LD's time for mountain
    • Babysit Mike Lesser's VPN, hopefully it's simple and will be done today

GE

  • Last week:
    • Scheduled the LUCI database cleanup with M.Smithwright and R.Reynolds for the week of July 25th, while LUCI-1 is off telescope.
    • Reviewed SX SOUL-LUCI handover documents, and added a couple of RIC's to address concerns. Observed SX SOUL-LUCI Verification Matrix testing.
    • Tested GEIRS double-buffering with D.Thompson to reduce overheads. In progress.
    • Investigated offsets and presets with OT group to resolve IT 8605. In progress.
    • Investigated alternative methods of launching the RTD from the Observer GUI. Found a way to activate a remote Java Console on the fly and updated the LUCI scripts.
  • This week:
    • I will be working half days (mornings 8-noon) most days through July 20th.
    • IT 8571, 8529 (Observer Scheduler GUI).

SH

  • troubles with UITS vis-a-vis Steward network
  • room 507 work
  • SHARK-NIR work and cleanup

next:

  • core switches
  • mountain trip

PSG

  • Last week
    • SHARK-NIR issues.
    • PDU is inaccessible - tried troubleshooting, called tech support.
    • FITS header information was duplicated, issue is fixed.
    • Simbad lookup fails, crashes process.
  • This week
    • Fix PDU for SHARK-NIR.
    • Fix Simbad lookup.
    • Continue addressing Issuetraks.

BM

  • Last week:
    • SX SOUL commissioning: more review of SW documents, watched tests
    • Update for AOIce patch (#108). Should be ready to checkout on restart.
    • More anemometer moxa connection investigating, updates to issue #144. (PR #143, test plan).
    • Reviewing old AO SW ITs
  • This week
    • CorrectModesAo error checking
    • RTDB disconnects (7213, 8397, 8628), anemometer disconnects (8241, 8285, 8341, 8492, 8656)
    • ITs (reviewing logs): 8307 (offsetxy), 8553/8250/3957 (bcu timeouts + ocam not live)
    • ATT meeting, SHARK-NIR Thursday re: taking images with soul wfs.
Topic revision: r8 - 11 Jul 2022, BrandonMechtley
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback